# CMDAR: A Chinese Multi-scene Dynamic Audio Reasoning Benchmark with Diverse Challenges

### v2

@@ 1 Introduction
The ability to reason from audio, including speech, paralinguistic cues, environmental sounds, and music, is essential for AI agents to interact effectively in real-world scenarios. Real-world environments are rich in overlapping auditory signals that convey linguistic content, social cues, temporal patterns, and environmental context (Wei et al., 2022; Nam, 2025). Effective reasoning over these audio streams is critical for AI agents to understand dynamic interactions, anticipate unfolding events, and support decision-making in applications such as autonomous navigation, audio event monitoring, and embodied household agents (Chen et al., 2020; Gao et al., 2023; Yeow et al., 2024).
@@
So far, various benchmarks have been proposed. AudioBench (Wang et al., 2025a) and AIR-Bench (Yang et al., 2024) are benchmark designed for large audio language models, aiming to comprehensively evaluate the understanding and reasoning ability of these models on audio content, but only focusing on shallow reasoning ability. MMAU (Sakshi et al., 2025) and MMAR focus more on the evaluation of deep reasoning ability of audio, but they pay more attention to perception samples, including counting and classification perception tasks, mainly for single scene and single audio problems, and do not include the evaluation of open-ended tasks, resulting in insufficient evaluation. Recently, MMAU-pro (Kumar et al., 2025) addressed the evaluation problem of multi-audio, but all the above evaluations only focus on English data, and do not focus on the investigation of Chinese audio reasoning ability and instruction following ability. To address the limitations of current audio reasoning benchmarks, we propose CMDAR, a Chinese comprehensive test benchmark aimed at evaluating the complex logical reasoning ability of LALMs in dynamic scenarios. CMDAR has 3000 high-quality questions from real scenarios, with multiple types of reasoning tasks (including continuous scenarios, causal scenarios, and multi scenario tasks) and comprehensive question types. All questions and reference answers are strictly reviewed by experts. The differences between the existing benchmark and CMDAR are summarized in Figure 1. In addition, we also propose a high-quality data construction process and a comprehensive and effective evaluation method.
@@
We benchmark 26 state-of-the-art audio language models on CMDAR and observe that they exhibit limitations in complex reasoning tasks. In the CMDAR-main, Qwen2.5-Omni (open-source) achieves 76.67% accuracy, whereas GPT-4o Audio (closed-source) reaches 68.47%. Among different tasks of single-choice questions, Qwen2.5-Omni (Xu et al., 2025a) performs worst on temporal reasoning (71.43%), while GPT-4o Audio struggles most on scene reasoning (61.27%). However, GPT-4o Audio substantially outperforms Qwen2.5-Omni on the more challenging multiple-choice with multiple audios and open-ended tasks. Across all three question types, no model surpasses 80% performance. These results indicate that even the most powerful models currently still have significant room for improvement, and finally we provide corresponding suggestions for the future development of LALMs. These findings underscore the unique challenges posed by CMDAR and its value as a benchmark for advancing audio reasoning research. Overall, our contributions are three-fold:
@@ 2.1 Audio-Language Models
Recent advances in large language models (LLMs) and cross-modal learning have driven substantial progress in audio-language research. Typical approaches jointly leverage speech, music, and general audio data, followed by instruction tuning to endow models with comprehension and instruction-following abilities. Open-source systems such as Qwen-Audio-Chat (Chu et al., 2023), Qwen2-Audio-Instruct (Chu et al., 2024), Audio Flamingo (Ghosh et al., 2025; Goel et al., 2025), SALMONN (Tang et al., 2024), and DeSTA2.5-Audio (Lu et al., 2025) report competitive results. More recent efforts, including Audio-CoT (Ma et al., 2025a), Audio-Reasoner (Xie et al., 2025), and R1-AQA (Li et al., 2025), focus on multi-step reasoning and challenging audio QA. Among proprietary systems, GPT-4o-Audio achieves the highest overall performance, while other commercial models such as Kimi-Audio (KimiTeam et al., 2025) and MiDashengLM (Dinkel et al., 2025) are also evaluated. Fully multimodal models such as Omni-R1 (Zhong et al., 2025) and Qwen2.5-Omni (Xu et al., 2025a) advance audio-text reasoning. We also benchmark cascaded audio-language models for a holistic comparison.
@@ 2.2 Audio Understanding and Reasoning Benchmarks
Prior works have explored audio question answering and compositional reasoning across speech, music, and environmental sounds. Early efforts include Lipping et al. (2022), a crowdsourced dataset with yes/no and single-word answers, and Ghosh et al. (2024b), which targets order- and attribute-level compositional reasoning with composition-aware fine-tuning for CLAP (Wu et al., 2023). In music, Melechovský et al. (2024) provides a controllable text-to-music system with theory-informed captions, and Weck et al. (2024) offers a human-validated multiple-choice benchmark probing music knowledge. For speech, Zhao et al. (2024) curates large-scale spoken QA in free-form and multiple-choice formats, while Huang et al. (2023) establishes a collaborative instruction-tuning benchmark across diverse speech tasks. Broader evaluations include Wang et al. (2025a), covering speech, scenes, and paralinguistics under instruction following, and Yang et al. (2024), which assesses generative comprehension and chat-based interaction over speech, sounds, and music. More recent reasoning-focused suites such as Sakshi et al. (2025), Ma et al. (2025b), and Wang et al. (2025b) emphasize multi-step perception and domain knowledge, revealing persistent gaps in multimodal integration and deep audio reasoning.
@@ 3.1 Overview of Benchmark
Audio Flamingo 2 (Ghosh et al., 2025) 3B LALMs 26.37 22.54 34.29 29.41 31.61 27.73
Audio Flamingo 3 (Goel et al., 2025) 7B LALMs 24.00 24.28 21.43 20.83 16.67 22.20
@@
DeSTA2.5-Audio (Lu et al., 2025) 8B LALMs 63.41 62.01 54.91 53.45 64.29 60.93
@@ 3.3 Evaluation metrics
CMDAR-multi. We use the following four metrics to measure model performance inspired by SATA-BENCH (Xu et al., 2025b), i) Exact Match (EM): For each sample, if the model’s predicted answer is as same as the correct answer, the accuracy is counted as 1; otherwise, it is counted as 0. The final EM is calculated as the average accuracy.
@@ 4.1 Setting
We evaluate the performance of both cascaded and non-cascaded models on CMDAR. The non-cascaded models include large audio-language models, large audio-reasoning models, and full-modal language models. The large audio-language models consist of Qwen-Audio-Chat (Chu et al., 2023), Qwen2-Audio-Instruct (Chu et al., 2024), SALAMONN (Tang et al., 2024), Audio Flamingo 2 (Ghosh et al., 2025), Audio Flamingo 3 Chat (Goel et al., 2025), MiDashengLM (Dinkel et al., 2025), etc. The large audio-reasoning models include Audio-Reasoner (Xie et al., 2025), etc. The full-modal language models include Omni-R1 (Zhong et al., 2025), Qwen2.5-Omni (Xu et al., 2025a), etc. The cascaded models based on captions include GPT-4o Audio + Qwen2-Audio-Instruct, Qwen2-Audio-Instruct + Llama-3-Instruct, etc. Considering the scarcity of Chinese audio reasoning data, some models lack of ability of Chinese reasoning or support for Chinese audio understanding. We annotate in the Table 1 whether they were trained using Chinese corpus to ensure fairness in evaluation. Please note that as our benchmark tests are automatically synthesized by Gemini, the results of Gemini related models will not be included. For the selection of models and detailed information, refer to the Appendix F.
@@ 4.3 Discussion
Gaussian Noise Replacement Experiment: Taking four different types of models as representatives, we replace the audio with Gaussian white noise of the same length and input it into the audio-language models. As shown in Figure 4, the accuracy of the models dropped significantly when they received white noise, approaching random guessing. This indicates that the models received audio information. However, the results also show that even with noisy audio, the models were still able to derive some answers from the textual questions.
@@ Appendix C Prompt of Evaluation
Prompt template for open-ended question answer scoring. System: "You are a helpful and precise assistant for checking the quality of the answer."
"[Detailed Audio Description][Audio]"
"[Question][Question]"
"[The Start of Assistant 1s Answer][Assistant1]"
"[The End of Assistant 1s Answer]"
"[The Start of Assistant 2s Answer][Assistant2]"
"[The End of Assistant 2s Answer]"
"[System]"
"We would like to request your feedback on the performance of two AI assistants in response to the user question "
"and audio description displayed above. AI assistants are provided with detailed audio descriptions and questions."
"Please rate the helpfulness, relevance, accuracy, and comprehensiveness of their responses. "
"Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance. "
"Please output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. "
"The two scores are separated by a space."
Assistant: {response}
Detailed explanation of the grading criteria. refer to AIR-Bench (Yang et al., 2024) System: "helpfulness": "Whether the response is helpful to the user and whether it can solve the user’s problem"
"relevance": "Whether the response is related to the question and audio content"
"accuracy": "The accuracy of the response, whether it is factual"
"comprehensiveness": "The comprehensiveness of the response, whether it covers all aspects of the question"
"score range": "1-10 points, the higher the score, the better the performance"
"output format": "Output a line containing two scores, separated by a space"
Assistant:
{response}
@@ Appendix D Prompt of Data construction pipeline
Prompt of Interference option generation System: You are a helpful AI assistant, designed to provide useful answer to customers. The user will provide you with a video, audio, audio translation, audio description, and video description. Please refer to this desired output and provide your own unique response. User: Based on the given QA pair (question and correct answer) and audio, video descriptions, and audio descriptions, generate 3 high-quality distractor options. the audio translation is [audio translation], the audio description is [audio description], and the video description is [video description]. The distractor options should possess deception (appearing reasonable from a surface logic or common sense perspective, easily causing those who lack key information to misselect), relevance (closely adhering to the question theme without deviating from the core scene/field), and incorrectness (clearly contradicting the correct answer or contradicting the facts in the background information), ultimately forming a complete set of multiple-choice options consisting of ’1 correct answer and 3 distractor options’. Assistant: {response}
@@ Appendix E Manual Screening Criteria
    Strictly check whether the question contains words such as "video" or "description".
@@ F.1 Evaluated Models
    Qwen2-Audio-Instruct  (Chu et al., 2024) is a new series of Qwen large audio-language models. It can accept inputs of various audio signals and perform audio analysis based on voice instructions or generate direct text responses. In this paper, we use the 7B Instruct model.
@@
    Qwen-Audio-Chat (Chu et al., 2023) is a Large Audio Language Model developed by Alibaba Cloud. It can take multiple types of audio (including human speech, natural sounds, music, and singing voices) and text as inputs, and generate text as output, supporting a multi-task learning framework for various complex audio tasks.
@@
    Audio Flamingo 2 (Ghosh et al., 2025) possesses advanced audio understanding and reasoning capabilities. In particular, it has professional audio reasoning ability and can understand long audio clips up to 5 minutes in duration.
@@
    Audio Flamingo 3 (Goel et al., 2025) (AF3) is a fully open-source, state-of-the-art Large Audio Language Model (LALM), which consists of an AF-Whisper unified audio encoder, an MLP-based audio adapter, a decoder-only LLM backbone (Qwen2.5-7B), and a streaming TTS module (AF3-Chat). Audio Flamingo 3 can accept audio inputs with a duration of up to 10 minutes.
@@
    Kimi-Audio-Instruct (KimiTeam et al., 2025) is designed as a general-purpose audio foundation model, which can handle a wide range of audio processing tasks within a single unified framework. It adopts mixed audio inputs (continuous acoustic signals + discrete semantic tokens) and an LLM core with parallel heads for text and audio token generation.
@@
    Omni-R1 (Zhong et al., 2025) is an all-modal model that addresses the resolution issue through a dual-system architecture. Meanwhile, it proposes an end-to-end RL framework—Omni-R1 is built on Group Relative Policy Optimization (GRPO). The results demonstrate the first successful application of RL in large-scale all-modal reasoning and highlight a scalable path toward general-purpose foundation models.
@@
    R1-AQA (Li et al., 2025) is an Audio Question Answering (AQA) model optimized through reinforcement learning using the Group Relative Policy Optimization (GRPO) algorithm.
@@
    SALAMONN (Tang et al., 2024) is a Large Language Model (LLM) that supports speech, audio event, and music inputs, developed by the Department of Electronic Engineering of Tsinghua University and ByteDance. It can perceive and understand various types of audio inputs. In this paper, we use the 7B version.
@@
    Audio-Reasoner (Xie et al., 2025) is an open-source project developed by a team from Tsinghua University, focusing on building a large language model that supports in-depth audio reasoning. Based on Qwen2-Audio-Instruct, this model incorporates structured chain-of-thought technology to achieve complex reasoning and multi-modal understanding of audio content.
@@
    DeSTA2.5-Audio (Lu et al., 2025) is a general-purpose Large Audio Language Model (LALM) designed for robust auditory perception and instruction-following capabilities, without the need for task-specific audio instruction tuning.
@@
    MiDashengLM (Dinkel et al., 2025) integrates the powerful Dasheng audio encoder with the Qwen2.5-Omni-7B Thinker decoder through a unique caption-based alignment strategy.
@@
    Qwen2.5-Omni (Xu et al., 2025a) is an end-to-end multi-modal model designed to perceive multiple modalities, including text, images, audio, and video, while generating text and natural speech responses in a streaming manner.

### v3

@@ 1 Introduction
The ability to reason from audio, including speech, paralinguistic cues, environmental sounds, and music, is essential for AI agents to interact effectively in real-world scenarios. Real-world environments are rich in overlapping auditory signals that convey linguistic content, social cues, temporal patterns, and environmental context (DBLP:journals/corr/abs-2208-09579; Auditory-Intelligence). Effective reasoning over these audio streams is critical for AI agents to understand dynamic interactions, anticipate unfolding events, and support decision-making in applications such as autonomous navigation, audio event monitoring, and embodied household agents (SoundSpaces; Sonicverse; Real-Time-Sound-Event).
@@
So far, various benchmarks have been proposed. AudioBench (AudioBench) and AIR-Bench (AIR-Bench) are benchmark designed for large audio language models, aiming to comprehensively evaluate the understanding and reasoning ability of these models on audio content, but only focusing on shallow reasoning ability. MMAU (MMAU) and MMAR focus more on the evaluation of deep reasoning ability of audio, but they pay more attention to perception samples, including counting and classification perception tasks, mainly for single scene and single audio problems, and do not include the evaluation of open-ended tasks, resulting in insufficient evaluation. Recently, MMAU-pro (MMAU-pro) addressed the evaluation problem of multi-audio, but all the above evaluations only focus on English data, and do not focus on the investigation of Chinese audio reasoning ability and instruction following ability. To address the limitations of current audio reasoning benchmarks, we propose CMDAR, a Chinese comprehensive test benchmark aimed at evaluating the complex logical reasoning ability of LALMs in dynamic scenarios. CMDAR has 3000 high-quality questions from real scenarios, with multiple types of reasoning tasks (including continuous scenarios, causal scenarios, and multi scenario tasks) and comprehensive question types. All questions and reference answers are strictly reviewed by experts. The differences between the existing benchmark and CMDAR are summarized in Figure 1. In addition, we also propose a high-quality data construction process and a comprehensive and effective evaluation method.
@@
We benchmark 26 state-of-the-art audio language models on CMDAR and observe that they exhibit limitations in complex reasoning tasks. In the CMDAR-main, Qwen2.5-Omni (open-source) achieves 76.67% accuracy, whereas GPT-4o Audio (closed-source) reaches 68.47%. Among different tasks of single-choice questions, Qwen2.5-Omni (Qwen2.5-Omni) performs worst on temporal reasoning (71.43%), while GPT-4o Audio struggles most on scene reasoning (61.27%). However, GPT-4o Audio substantially outperforms Qwen2.5-Omni on the more challenging multiple-choice with multiple audios and open-ended tasks. Across all three question types, no model surpasses 80% performance. These results indicate that even the most powerful models currently still have significant room for improvement, and finally we provide corresponding suggestions for the future development of LALMs. These findings underscore the unique challenges posed by CMDAR and its value as a benchmark for advancing audio reasoning research. Overall, our contributions are three-fold:
@@ 2.1 Audio-Language Models
Recent advances in large language models (LLMs) and cross-modal learning have driven substantial progress in audio-language research. Typical approaches jointly leverage speech, music, and general audio data, followed by instruction tuning to endow models with comprehension and instruction-following abilities. Open-source systems such as Qwen-Audio-Chat (Qwen-Audio-Chat), Qwen2-Audio-Instruct (Qwen2-Audio-Instruct), Audio Flamingo (Audio-Flamingo-2; Audio-Flamingo-3), SALMONN (SALMONN), and DeSTA2.5-Audio (DeSTA2.5-Audio) report competitive results. More recent efforts, including Audio-CoT (Audio-CoT), Audio-Reasoner (Audio-Reasoner), and R1-AQA (R1-AQA), focus on multi-step reasoning and challenging audio QA. Among proprietary systems, GPT-4o-Audio achieves the highest overall performance, while other commercial models such as Kimi-Audio (Kimi-audio) and MiDashengLM (Midashenglm) are also evaluated. Fully multimodal models such as Omni-R1 (Omni-R1) and Qwen2.5-Omni (Qwen2.5-Omni) advance audio-text reasoning. We also benchmark cascaded audio-language models for a holistic comparison.
@@ 2.2 Audio Understanding and Reasoning Benchmarks
Prior works have explored audio question answering and compositional reasoning across speech, music, and environmental sounds. Early efforts include Clotho-AQA, a crowdsourced dataset with yes/no and single-word answers, and CompA, which targets order- and attribute-level compositional reasoning with composition-aware fine-tuning for CLAP (CLAP). In music, Mustango provides a controllable text-to-music system with theory-informed captions, and MuChoMusic offers a human-validated multiple-choice benchmark probing music knowledge. For speech, LibriSQA curates large-scale spoken QA in free-form and multiple-choice formats, while Dynamic-SUPERB establishes a collaborative instruction-tuning benchmark across diverse speech tasks. Broader evaluations include AudioBench, covering speech, scenes, and paralinguistics under instruction following, and AIR-Bench, which assesses generative comprehension and chat-based interaction over speech, sounds, and music. More recent reasoning-focused suites such as MMAU, MMAR, and MMSU emphasize multi-step perception and domain knowledge, revealing persistent gaps in multimodal integration and deep audio reasoning.
@@ 3.1 Overview of Benchmark
Audio Flamingo 2 (Audio-Flamingo-2) 3B LALMs 26.37 22.54 34.29 29.41 31.61 27.73
Audio Flamingo 3 (Audio-Flamingo-3) 7B LALMs 24.00 24.28 21.43 20.83 16.67 22.20
@@
DeSTA2.5-Audio (DeSTA2.5-Audio) 8B LALMs 63.41 62.01 54.91 53.45 64.29 60.93
@@ 3.3 Evaluation metrics
CMDAR-multi. We use the following four metrics to measure model performance inspired by SATA-BENCH (SATA), i) Exact Match (EM): For each sample, if the model’s predicted answer is as same as the correct answer, the accuracy is counted as 1; otherwise, it is counted as 0. The final EM is calculated as the average accuracy.
@@ 4.1 Setting
We evaluate the performance of both cascaded and non-cascaded models on CMDAR. The non-cascaded models include large audio-language models, large audio-reasoning models, and full-modal language models. The large audio-language models consist of Qwen-Audio-Chat (Qwen-Audio-Chat), Qwen2-Audio-Instruct (Qwen2-Audio-Instruct), SALAMONN (SALMONN), Audio Flamingo 2 (Audio-Flamingo-2), Audio Flamingo 3 Chat (Audio-Flamingo-3), MiDashengLM (Midashenglm), etc. The large audio-reasoning models include Audio-Reasoner (Audio-Reasoner), etc. The full-modal language models include Omni-R1 (Omni-R1), Qwen2.5-Omni (Qwen2.5-Omni), etc. The cascaded models based on captions include GPT-4o Audio + Qwen2-Audio-Instruct, Qwen2-Audio-Instruct + Llama-3-Instruct, etc. Considering the scarcity of Chinese audio reasoning data, some models lack of ability of Chinese reasoning or support for Chinese audio understanding. We annotate in the Table 1 whether they were trained using Chinese corpus to ensure fairness in evaluation. Please note that as our benchmark tests are automatically synthesized by Gemini, the results of Gemini related models will not be included. For the selection of models and detailed information, refer to the Appendix F.
@@ 4.3 Discussion
Gaussian Noise Replacement Experiment: Taking four different types of models as representatives, we replace the audio with Gaussian white noise of the same length and input it into the audio-language models. As shown in the middle of Figure 4, the accuracy of the models dropped significantly when they received white noise, approaching random guessing. This indicates that the models received audio information. However, the results also show that even with noisy audio, the models were still able to derive some answers from the textual questions.
@@ Appendix C Prompt of Evaluation
Prompt template for open-ended question answer scoring. System: “You are a helpful and precise assistant for checking the quality of the answer.” “[Detailed Audio Description][Audio]” “[Question][Question]” “[The Start of Assistant 1s Answer][Assistant1]” “[The End of Assistant 1s Answer]” “[The Start of Assistant 2s Answer][Assistant2]” “[The End of Assistant 2s Answer]” “[System]” “We would like to request your feedback on the performance of two AI assistants in response to the user question ” “and audio description displayed above. AI assistants are provided with detailed audio descriptions and questions.” “Please rate the helpfulness, relevance, accuracy, and comprehensiveness of their responses. ” “Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance. ” “Please output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. ” “The two scores are separated by a space.” Assistant: {response}
Detailed explanation of the grading criteria. refer to AIR-Bench (AIR-Bench) System: “helpfulness”: “Whether the response is helpful to the user and whether it can solve the user’s problem” “relevance”: “Whether the response is related to the question and audio content” “accuracy”: “The accuracy of the response, whether it is factual” “comprehensiveness”: “The comprehensiveness of the response, whether it covers all aspects of the question” “score range”: “1-10 points, the higher the score, the better the performance” “output format”: “Output a line containing two scores, separated by a space” Assistant: {response}
@@ Appendix D Prompt of Data construction pipeline
Prompt of Interference option generation System: You are a helpful AI assistant, designed to provide useful answer to customers. The user will provide you with a video, audio, audio translation, audio description, and video description. Please refer to this desired output and provide your own unique response. User: Based on the given QA pair (question and correct answer) and audio, video descriptions, and audio descriptions, generate 3 high-quality distractor options. the audio translation is [audio translation], the audio description is [audio description], and the video description is [video description]. The distractor options should possess deception (appearing reasonable from a surface logic or common sense perspective, easily causing those who lack key information to misselect), relevance (closely adhering to the question theme without deviating from the core scene/field), and incorrectness (clearly contradicting the correct answer or contradicting the facts in the background information), ultimately forming a complete set of multiple-choice options consisting of “1 correct answer and 3 distractor options”. Assistant: {response}
@@ Appendix E Manual Screening Criteria
    Strictly check whether the question contains words such as “video” or “description”.
@@ F.1 Evaluated Models
    Qwen2-Audio-Instruct  (Qwen2-Audio-Instruct) is a new series of Qwen large audio-language models. It can accept inputs of various audio signals and perform audio analysis based on voice instructions or generate direct text responses. In this paper, we use the 7B Instruct model.
@@
    Qwen-Audio-Chat (Qwen-Audio-Chat) is a Large Audio Language Model developed by Alibaba Cloud. It can take multiple types of audio (including human speech, natural sounds, music, and singing voices) and text as inputs, and generate text as output, supporting a multi-task learning framework for various complex audio tasks.
@@
    Audio Flamingo 2 (Audio-Flamingo-2) possesses advanced audio understanding and reasoning capabilities. In particular, it has professional audio reasoning ability and can understand long audio clips up to 5 minutes in duration.
@@
    Audio Flamingo 3 (Audio-Flamingo-3) (AF3) is a fully open-source, state-of-the-art Large Audio Language Model (LALM), which consists of an AF-Whisper unified audio encoder, an MLP-based audio adapter, a decoder-only LLM backbone (Qwen2.5-7B), and a streaming TTS module (AF3-Chat). Audio Flamingo 3 can accept audio inputs with a duration of up to 10 minutes.
@@
    Kimi-Audio-Instruct (Kimi-audio) is designed as a general-purpose audio foundation model, which can handle a wide range of audio processing tasks within a single unified framework. It adopts mixed audio inputs (continuous acoustic signals + discrete semantic tokens) and an LLM core with parallel heads for text and audio token generation.
@@
    Omni-R1 (Omni-R1) is an all-modal model that addresses the resolution issue through a dual-system architecture. Meanwhile, it proposes an end-to-end RL framework—Omni-R1 is built on Group Relative Policy Optimization (GRPO). The results demonstrate the first successful application of RL in large-scale all-modal reasoning and highlight a scalable path toward general-purpose foundation models.
@@
    R1-AQA (R1-AQA) is an Audio Question Answering (AQA) model optimized through reinforcement learning using the Group Relative Policy Optimization (GRPO) algorithm.
@@
    SALAMONN (SALMONN) is a Large Language Model (LLM) that supports speech, audio event, and music inputs, developed by the Department of Electronic Engineering of Tsinghua University and ByteDance. It can perceive and understand various types of audio inputs. In this paper, we use the 7B version.
@@
    Audio-Reasoner (Audio-Reasoner) is an open-source project developed by a team from Tsinghua University, focusing on building a large language model that supports in-depth audio reasoning. Based on Qwen2-Audio-Instruct, this model incorporates structured chain-of-thought technology to achieve complex reasoning and multi-modal understanding of audio content.
@@
    DeSTA2.5-Audio (DeSTA2.5-Audio) is a general-purpose Large Audio Language Model (LALM) designed for robust auditory perception and instruction-following capabilities, without the need for task-specific audio instruction tuning.
@@
    MiDashengLM (Midashenglm) integrates the powerful Dasheng audio encoder with the Qwen2.5-Omni-7B Thinker decoder through a unique caption-based alignment strategy.
@@
    Qwen2.5-Omni (Qwen2.5-Omni) is an end-to-end multi-modal model designed to perceive multiple modalities, including text, images, audio, and video, while generating text and natural speech responses in a streaming manner.